OpenStudy (anonymous):

help fast! Find the 5th term of the geometric sequence 9, 27, 81,…

OpenStudy (anonymous):

GS terms are a, ar, ar^2 ... a = first term, r = common ratio

OpenStudy (campbell_st):

use the formula \[T_{n} = ar^{n -1}\] a = 9 and r = 3 n = 5

OpenStudy (anonymous):

i tried, what would the fifth term be?

OpenStudy (anonymous):

9 * 3^4

OpenStudy (campbell_st):

\[T_{5} = 9\times3^{5 -1}\] so \[T_{5} = 9 \times 3^4\] here is a calculator you can use http://web2.0calc.com/

OpenStudy (anonymous):

9* 3^4 can u just do that? trust me i know how to tho

OpenStudy (campbell_st):

well its 9 x 3 x 3 x 3 x 3= 9 x 81

OpenStudy (anonymous):

81x9?

OpenStudy (campbell_st):

yep that 9 x 80 + 9 x 1 =

OpenStudy (anonymous):

279??

OpenStudy (anonymous):

is that my answer?

OpenStudy (campbell_st):

close... swap the first 2 digits

OpenStudy (anonymous):

729 oops!

OpenStudy (campbell_st):

thats the answer